Our environment is very sensitive to the amount of ozone in the upper atmosphere. The level of ozone normally found is 5.7 parts/million (ppm). A researcher believes that the current ozone level is at an excess level. The mean of 10 samples is 6.1 ppm with a variance of 0.25. Does the data support the claim at the 0.01 level? Assume the population distribution is approximately normal. Step 4 of 5 : Determine the decision rule for rejecting the null hypothesis. Round your answer to three decimal places.
Answer #1

H0: Null Hypothesis: \mu = 5.7

HA: Alternative Hypothesis: \mu > 5.7

SE = s/\sqrt{n}

= \sqrt{25}/\sqrt{10}

= 1.5811

Test Statistic is given by:

t = (6.1 - 5.7)/1.5811

= 0.253

\alpha = 0.01

ndf = n - 1 = 10 - 1 = 9

One Tail - Right Side Test

From Table, critical value of t = 2.821

Decision Rule:
Reject null hypothesis if t > 2.821

Since the calculated value of t is less than critical value, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data do not support the claim that the current ozone level is at an excess level.

Answer to Question asked:

Question:

Determine the decision rule for rejecting the null hypothesis.

Answer :

Reject null hypothesis if t > 2.821